# Qubit-Efficient Quantum Annealing for Stochastic Unit Commitment

### v1

@@ Abstract
Stochastic Unit Commitment (SUC) has been proposed to manage the uncertainties driven by the integration of renewable energy sources. When solved by Benders Decomposition (BD), the master problem becomes a binary integer programming which is NP-hard and computationally demanding for classical computational methods. Quantum Annealing (QA), known for efficiently solving Quadratic Unconstrained Binary Optimization (QUBO) problems, presents a potential solution. However, existing quantum algorithms rely on slack variables to handle linear binary inequality constraints, leading to increased qubit consumption and reduced computational efficiency. To solve the problem, this paper introduces the Powell-Hestenes-Rockafellar Augmented Lagrangian Multiplier (PHR-ALM) method to eliminate the need for slack variables so that the qubit consumption becomes independent of the increasing number of bender’s cuts. To further reduce the qubit overhead, quantum ADMM is applied to break large-scale SUC into smaller blocks and enables a sequential solution. Consequently, the Quantum-based PHR-ADMM (QPHR-ADMM) can significantly reduce qubit requirements and enhancing the applicability of QA in SUC problem. The simulation results demonstrate the feasibility of the proposed QPHR-ADMM algorithm, indicating its superior time efficiency over classical approaches for large scale QUBO problems under the D-Wave QPU showcases.
Index Terms:
Quantum computing, quantum annealing, two-stage stochastic unit commitment programming, benders decomposition, PHR-Augmented Lagrangian Multiplier, D-ADMM.
@@ I Introduction
AS the penetration of renewable energy, such as wind and solar, continues to increase within power systems, how to efficiently handle the uncertainties associated with these resources in system operation remains an open question [1].
@@ I-A Unit Commitment Problem
To ensure secure and reliable power system operation, the system operator implements unit commitment (UC) in day ahead to decide the online status of the generators that minimizes production costs while ensuring supply and demand balance, given energy forecasts and various operational constraints. The problem can be mathematically formulated as a mixed-integer programming (MIP) problem [2], which is NP-hard [3]. The complexity and computational burden of UC poses significant challenges to classic computational methods.
Depending on different treatments of uncertainties, UC can be classified into deterministic, robust, and stochastic UC. In deterministic UC models, parameters such as load demand are assumed to be precisely known, excluding the representation of uncertainty and resulting in a relatively smaller problem scale. Unlike deterministic models, the robust [4, 5, 6] and stochastic [7, 8, 9] UC problems offer a more detailed and realistic representation of power system operations by explicitly accounting for uncertainties through the introduction of random variables [10]. In SUC, uncertainty is modeled using probabilistic methods by generating a representative set of scenarios through sampling. However, achieving a reliable and feasible solution may require a large number of scenarios, significantly increasing computational complexity. Multi-block Alternating Direction Method of Multipliers (ADMM) and Benders Decomposition (BD) are two algorithms that can efficiently reduce the problem size and allow parallel processing. However, dealing with large numbers of integer variables and scenarios is still challenging for a real-time application.
@@ I-B Quantum Computing for Power System Operations
Quantum computing (QC), an emerging computational paradigm grounded in quantum mechanics[11], provides a novel architecture with the potential to overcome current computational bottlenecks. Unlike classical computing, where the fundamental unit is the bit, QC uses quantum bits, or qubits, to store information. The essence of QC involves exploiting quantum entanglement and superposition to create and manipulate complex multi-eigenstate superpositions [11]. This capability allows quantum computers to process multiple eigenstates in parallel. Currently, QC has been developed under two primary computational models: the gate-based model and the quantum annealing (QA) based model [12]. The gate-based model constructs quantum circuits using quantum gates to manipulate qubit states, thereby evolving the probabilities of all eigenstates to achieve specific computational goals. On the other hand, QA utilizes the phenomenon of quantum tunneling to identify the lowest energy state of the system, known as the ground state [13]. In this ground state, the spin orientations of qubits can be interpreted as the global optimal solution to a Quadratic Unconstrained Binary Optimization (QUBO).
@@
In power systems, the potential of QC algorithms has already been explored in initial applications. For example, the Harrow-Hassidim-Lloyd (HHL) algorithm provides exponential acceleration to solve linear problems, including DC optimal power flow (DCOPF) [14] and fast decoupled load flow (FDLF) [15], compared to classical algorithms. [16] introduces a Quantum Monte Carlo Simulation (MCS) method based on Quantum Amplitude Estimation (QAE) for power system reliability assessment under uncertainty, demonstrating a quadratic speedup over classical MCS. Moreover, Quantum Approximate Optimization Algorithm (QAOA) is a promising method for realizing quantum speedup in solving combinatorial optimization problems. For instance, QAOA has been used to determine optimal planning strategies for efficient electric vehicle charging within the power grid [17]. Furthermore, QA has been applied to solve combinatorial optimal power flow problems using the D-Wave quantum processing unit (QPU), as proposed in [18].
Beyond the applications mentioned above, QC offers emerging opportunities for UC. The UC problem can be reformulated into a QUBO format, enabling evolution through a quantum gate circuit designed by QAOA, leveraging Hamiltonian simulation [19, 20, 21]. Due to the limited capability of QC in handling continuous variables, classical-quantum hybrid methods have been proposed. For example, [20] introduces the Quantum ADMM (QADMM) algorithm to partition integer and continuous blocks. Similarly, [21] employs Quantum Surrogate Lagrangian Relaxation (QSLR) to coordinate quantum and classical resources locally and across neighboring subproblems for computation.
@@
Since the master problem in the BD framework encompasses all first-stage integer variables, QC is particularly suitable for SUC problems. A key challenge is how to efficiently manage the increasing number of Benders cut constraints during Bender’s iteration. Slack variables are often used to convert the resulting inequality constraints into equalities [21, 22, 23, 24, 25]. However, enabling QPUs to handle these slack variables requires a substantial allocation of qubits for binary encoding (a technique to convert continuous variables into binary variables). With each BD iteration building a new Benders cut, the inclusion of slack variables dramatically extends the qubit demand. This leads to significant qubit overhead, reduced computational efficiency, or even causes chain break, a catastrophic failure of QC.
@@ I-C Contributions
This paper, for the first time, studies the efficient use of QA for solving SUC within the BD framework. Combining the Powell-Hestenes-Rockafellar Augmented Lagrangian Multiplier (PHR-ALM) and Direct-extended ADMM (D-ADMM), Quantum-based PHR ADMM (QPHR-ADMM) is proposed to significantly reduce the qubit overhead, improve the QC accuracy, and enhance computational efficiency. The contributions of this paper can be summarized as:
@@
    Compared with conventional QC algorithms that handle inequality constraints with binary-encoded slack variables, the proposed PHR-ALM method eliminates slack variables within the BD framework. We theoretically demonstrate that this method breaks down the linear dependency between the number of qubits and the number of BD iterations.
@@
    By integrating with D-ADMM algorithm, QPHR-ADMM can break large-scale SUC into smaller blocks, which further reduces qubit numbers and enables a sequential solution. We theoretically prove that the proposed algorithm reduces the multiple linear space complexities into a constant which equals to the time horizon of the SUC or the pre-defined precision in binary encoding.
@@
    For both PHR-ALM and D-ADMM, we provide an analytical derivation of their QUBO and Hamiltonian representations, making them directly implementable with QA algorithms. Through a case study on SUC, we demonstrate that the proposed approach eliminates the need for auxiliary qubits on QPUs compared to basic QA while achieving superior qubit efficiency.
With advancements in quantum technology, QA is expected to surpass classical computation for solving SUC. More broadly, we demonstrate that even with limited quantum resources, the potential of QC can be effectively harnessed through rigorous mathematical reformulations, paving the way for quantum-integrated optimization for power system applications.
The structure of this paper is organized as follows. Section II describes the scenario-based SUC problem, providing its mathematical formulation within the BD framework, along with the overview of QC technology. Section III details the design process of the QPHR-ADMM algorithm. Two specific case studies are presented in Section IV to demonstrate the feasibility, accuracy, and efficiency of the proposed QPHR-ADMM algorithm. Finally, Section V concludes the paper.
@@ II Preliminaries
This section provides the background of SUC and explains how QA can be applied to solve integer optimization problems.
@@ II-A1 SUC Problem Formulation
To model the uncertainty of wind power generation on the supply side and the uncertainty of electricity consumption on the demand side, a two-stage scenario-based SUC model can be applied. The first stage involves making day-ahead UC decisions that satisfy all unit constraints, primarily determining the on/off status of thermal generators. The second stage involves the economic dispatch of thermal generators under each uncertain scenario, while ensuring compliance with operation constraints.
In detail, during an operation horizon of $T$ periods, the decision variables are categorized into two groups. The first-stage decision variables are binary variables $u_{{g},{t}}$ determining the on/off states for the $g$-th unit in time period $t$. Since these first-stage decisions pertain to day-ahead scheduling, they are independent of realizations of uncertainties. Subsequently, when wait-and-see decisions are made, uncertainties in the power system are realized in terms of the output of wind farms on the supply side and the load on the demand side. The corresponding random variables are defined as $P_{{h},{t}}^{Wind}$ and $D_{{h},{t}}^{Load}$, respectively, for period $t$ for scenario $h$ in the scenario set $\Omega$ with probability $\pi\left(\xi_{h}\right)$. Due to uncertainties in both supply and demand in the second stage, we introduce load shedding, modeled as a high-cost, fast-start, and unconstrained generator, to maintain the supply-demand balance, thereby enhancing feasibility. Therefore, the decision variables at this stage include not only the output of each unit in different time periods $P_{{h},{g},{t}}^{G}$ but also the load shedding amount $P_{{h},{t}}^{shed}$ during the corresponding periods.
A complete formulation of the scenario-based SUC is considered as follows[26, 27, 28].
@@
$\sum_{g=1}^{N}P_{h,g,t}^{G}+P_{h,t}^{shed}+P_{h,t}^{Wind}=D_{h,t}^{Load}\quad\forall h,g,t$ (1d)
$u_{g,t}\in\{0,1\}\quad\forall g,t$ (1e)
$P_{h,t}^{shed}\geq 0\quad\forall h,t$ (1f)
where $C_{g}^{quad}$, $C_{g}^{prim}$ and $C_{g}^{cons}$ are cost coefficients. $C^{shed}$ is the penalty cost of load shedding. $P_{g}^{min}$ and $P_{g}^{max}$ donate the minimum and maximum power output. $P_{g}^{rd}$ and $P_{g}^{ru}$ are ramp-up and ramp-down rate limits. With consideration of economic dispatches in SUC, the objective function aims to minimize the total cost, consisting of constant production costs and the quadratic fuel costs of thermal units in (1a). (1b) and (1c) are generator capacity and ramp limits, respectively. (1d) represents the power balance constraints for any given period in any scenario.
@@ II-A2 Solution Algorithm Based on Benders Decomposition
The BD method has been widely applied in large-scale power system optimization to reduce its computational burden[23, 24, 29]. This algorithm decouples complex, multi-constraint problems into small-scale subproblems, that can be potentially solved in parallel. In (1b), the first-stage variables are incorporated into the second-stage constraints, leading to a coupling between the two stages. To decouple the two stages, BD divides the SUC problem (1a)-(1f) into two sections, i.e., a mixed-integer master problem and several scenario-based quadratic subproblems. The master problem represents the first-stage commitment problem, where the variables only include the binary on/off statuses of the units $u_{g,t}^{k}$.
In detail, the master problem is written as,
$\min_{u_{g,t}^{k},\Upsilon^{k}}\limits\sum_{g=1}^{N}\limits\sum_{t=1}^{T}C_{g}^{cons}u_{g,t}^{k}+\Upsilon^{k}$ (2a)
@@
$\quad\!\!\left.+C_{g}^{prim}P_{h,g,t}^{G,l}\bigg{]}+\sum_{t=1}^{T}C^{shed}P_{h,t}^{shed,l}\right\}$
@@
$u_{g,t}^{k}\in\{0,1\}\quad\forall g,t$ (2d)
where $\theta_{g,t}^{l}\left(\xi_{h}\right)$s are the dual variables introduced by the subproblems. Note that, due to the introduction of a continuous Benders lower bound (LB), $\Upsilon$, the master problem of the SUC in the $k$-th iteration, given in (2a)-(2d), is characterized as an MIP problem rather than an integer programming problem.
As each scenario is independent, the subproblem is decoupled and can be treated separately. For example, the subproblem for scenario $h$ is given as (3a)-(3f) below. The main objective is to strategize the optimal power generation levels for each unit in different scenarios, along with the corresponding load shedding amount.
@@
In each subproblem, $u_{g,t}^{fixed}$ is a newly introduced auxiliary continuous variable and $u_{g,t}^{k}$ is the optimal decision variable of the master problem [2]. As a result, (3) results in an upper bound (UB) of (1). Meanwhile, due to the convexity of (3), the dual variable $\theta_{g,t}^{l}\left(\xi_{h}\right)$s in (3e) defines the sensitivity of subproblem’s objective. After collecting all sensitivities, a new Benders cut (2c) can be added in the next iteration and the master problem (2) becomes a LB to (1). Moreover, as subproblem is a continuous quadratic programming, it can be efficiently solved on CPUs.
The BD method decouples the first-stage and second-stage operation, significantly reducing the problem size. It is noteworthy that all integer variables are now included in the master problem where QA holds the potential to further improve its computation efficiency.

### v2

@@ Abstract
Stochastic Unit Commitment (SUC) has been proposed to manage the uncertainties driven by renewable integration, but it leads to significant computational complexity. When accelerated by Benders Decomposition (BD), the master problem becomes binary integer programming, which is still NP-hard and computationally demanding for classical methods. Quantum Annealing (QA), known for efficiently solving Quadratic Unconstrained Binary Optimization (QUBO) problems, presents a potential solution. However, existing quantum algorithms rely on slack variables to handle linear binary inequality constraints, leading to increased qubit consumption and reduced computational efficiency. To solve the problem, this paper introduces the Powell-Hestenes-Rockafellar Augmented Lagrangian Multiplier (PHR-ALM) method to eliminate the need for slack variables, making qubit consumption independent of the increasing number of Benders cuts. To further reduce the qubit overhead, quantum ADMM is applied to break large-scale SUC into smaller blocks for sequential solutions, which does not scale with the number of generators. Finally, the simulation results on both 4-generator and the IEEE bus-118 systems demonstrate the feasibility and scalability of the proposed algorithm, indicating its superior qubit and runtime efficiency over classical and baseline quantum approaches on the D-Wave QPU platform.
@@ I-A Unit Commitment Problem
AS the penetration of renewable energy continues to increase, how to efficiently handle the uncertainties associated with these resources in the system operation remains an open question [5]. To ensure secure and reliable operation, the system operator implements unit commitment (UC) in day ahead to decide the online status of the generators that minimizes production costs while ensuring supply and demand balance, given energy forecasts and various operational constraints. The problem can be mathematically formulated as a mixed-integer programming (MIP) problem [30], which is NP-hard [8] and poses significant challenges to classic computational methods.
Depending on different treatments of uncertainties, UC can be classified into deterministic, robust, and stochastic UC. In deterministic UC models, parameters such as load demand and renewable generation are assumed to be precisely known, resulting in a relatively smaller scale problem. Unlike deterministic models, the robust [44, 51, 25] and stochastic [52, 19, 3] UC problems explicitly account for uncertainties as random variables [42]. In SUC, uncertainty is modeled using probabilistic methods by generating a representative set of scenarios through sampling. However, achieving a representative and feasible solution may require a large number of scenarios, significantly increasing computational complexity. Multi-block Alternating Direction Method of Multipliers (ADMM) and Benders Decomposition (BD) are two algorithms that can efficiently reduce the problem size and allow parallel processing. However, dealing with large numbers of integer variables and scenarios is still challenging for a real-time application.
@@ I-B Quantum Computing for Power System Operations
Quantum computing (QC), an emerging computational paradigm grounded in quantum mechanics [31], provides a novel architecture with the potential to overcome current computational bottlenecks. Unlike classical computing, where the fundamental unit is the bit, QC uses quantum bits, or qubits, to store information. The essence of QC involves exploiting quantum entanglement and superposition to create and manipulate complex multi-eigenstate superpositions [31]. This capability allows quantum computers to process multiple eigenstates in parallel. Currently, QC has been developed under two primary computational models: the gate-based model and the quantum annealing (QA) based model [28]. The gate-based model constructs quantum circuits, such as quantum approximate optimization algorithm (QAOQ), uses quantum gates to manipulate qubit states, thereby evolving the probabilities of all eigenstates to achieve specific computational goals. On the other hand, QA utilizes the phenomenon of quantum tunneling to identify the lowest energy state of the system, known as the ground state [34]. In this ground state, the spin orientations of qubits can be interpreted as the global optimal solution to a Quadratic Unconstrained Binary Optimization (QUBO).
@@
In power systems, the potential of QC has been explored across problems of increasing structural complexity. Early efforts focused on linear system applications, where the Harrow-Hassidim-Lloyd algorithm provides exponential acceleration for solving linear problems such as DC optimal power flow [2] and fast decoupled load flow [11], compared to classical methods. Beyond deterministic linear models, QC has been extended to uncertainty quantification. For example, [33] proposes a Quantum Monte Carlo Simulation (MCS) method based on Quantum Amplitude Estimation for reliability assessment under uncertainty, demonstrating quadratic speedup over classical MCS.
More recently, QC has been applied to mixed-integer optimization problems in power systems, many of which can be reformulated into QUBO form. For instance, QAOA has been used for optimal electric vehicle charging [20], and QA has been employed to solve combinatorial optimal power flow [29] problems using D-Wave quantum processors. Building upon these developments, QC offers opportunities for more complex mixed-integer problems such as unit commitment (UC)[26, 32, 10, 16].
@@
Given that QC primarily targets integer programming, the BD framework is appealing because the master problem explicitly includes all first-stage integer variables in a two-stage problem such as SUC. Table I provides a comparison of recent BD-based quantum-assisted formulations for power system applications, highlighting their treatment of slack variables, additional qubit requirements, and convergence behavior. A key challenge lies in the efficient management of the increasing number of Benders cut constraints as the algorithm iterates. In detail, slack variables are often used to convert the resulting inequality constraints into equalities [26, 32, 10, 45, 23, 12, 16, 4, 35, 24, 13]. However, enabling QPUs to handle these slack variables requires a substantial allocation of qubits for binary encoding (a technique to convert continuous variables into binary variables), whose qubit occupation increases linearly with the Benders’ iterations. This leads to significant qubit overhead, reduced computational efficiency, or even causes chain break, a catastrophic failure mode of QC.
TABLE I: Recent BD-based quantum-assisted approaches for power system applications.
Reference Quantum Algorithm Problem Binary-encoded Slack Variables Source of Additional Qubits Component(s) with Convergence Guaranteed
Benders Cuts Power System Constraints
[23] QA Multi-energy System
[12] QA Post-disaster Restoration
[54] QA Energy Management
[48] QA Federated Learning Scheduling BD
[4] QA UC
[35] QA UC BD
[24] QA UC QA
[13] QAOA/QA Security-constrained UC BD and QAOA
This work QA SUC BD and QA
@@ I-C Contributions
This paper studies the efficient use of QA for solving SUC within the BD framework. Combining the Powell-Hestenes-Rockafellar Augmented Lagrangian Multiplier (PHR-ALM) and Direct-extended ADMM (D-ADMM), Quantum-based PHR ADMM (QPHR-ADMM) is proposed. The contributions of this paper are summarized as:
@@
    The PHR-ALM method is proposed to eliminate the need of slack variables for inequality constraint embedding within the BD framework, which breaks down the linear dependency between the number of qubits and the number of BD iterations.
@@
    D-ADMM decomposes large-scale SUC into smaller blocks, enabling sequential solves and eliminating the linear scaling with the number of generators; the resulting complexity depends primarily on the time horizon and the binary-encoding precision.
@@
    For both PHR-ALM and D-ADMM, we provide analytical derivations of their QUBO and Hamiltonian representations, making them directly implementable with QA algorithms.
  4. 4.
    This work extends the previous application of QA for UC problem. Notably, Prior works [16, 35, 24] on D-Wave QPUs are usually restricted to systems with maximum 15 units and often do not consider a full 24-hour scheduling horizon. By contrast, through various case studies on synthetic, 4-generator, and realistic IEEE bus-118 test cases, we demonstrate the qubit and computational time efficiencies as well as the scalability of the proposed algorithm, compared to baseline quantum methods.
To sum up, with advancements in quantum technology, QA is expected to surpass classical computation for solving SUC. More broadly, this paper demonstrates that even with limited quantum resources, the potential of QC can be effectively harnessed through rigorous mathematical reformulations, accelerating the deployment of quantum-integrated optimization for power system applications. The structure of this paper is organized as follows. Section II describes the scenario-based SUC problem, solved by BD framework, along with the overview of QC technology. Particularly, we discuss why QA is chosen, compared to gate-based method, such as QAOA. Section III details the design process of the QPHR-ADMM algorithm. Case studies are presented in Section IV. Finally, Section V concludes the paper.
@@ II-A1 SUC Problem Formulation
To model the uncertainty of wind power generation on the supply side and the uncertainty of electricity consumption on the demand side, a two-stage scenario-based SUC model can be applied. The first stage involves making day-ahead UC decisions that satisfy all unit constraints, primarily determining the on/off status of thermal generators. The second stage involves the economic dispatch of thermal generators under each uncertain scenario, while ensuring compliance with operation constraints. A complete formulation of the scenario-based SUC is considered as follows[41, 6, 43].
@@
$\sum_{g=1}^{N}P_{h,g,t}^{G}+P_{h,t}^{shed}+P_{h,t}^{Wind}=D_{h,t}^{Load}\quad\forall h,t$ (1d)
$\sum_{\varsigma=t}^{t+T_{g}^{U}-1}\!\!\!\!u_{g,\varsigma}\!\!\;\geq\;\!\!T_{g}^{U}\,\!\!\bigl(u_{g,t}\!-\!u_{g,t-1}\bigr)\!\!\!\quad\forall g,\;\!\!t\!=\!2,\!\dots\!,T\!-\!T_{g}^{U}\!\!+1\!\!\!$ (1e)
$\sum_{\varsigma=t}^{t+T_{g}^{D}-1}\!\!\!\!\!\bigl(1\!-\!u_{g,\varsigma}\bigr)\;\!\!\geq\;\!\!T_{g}^{D}\,\!\!\bigl(u_{g,t-1}\!-\!u_{g,t}\bigr)\!\!\!\!\quad\forall g,\;\!\!t\!=\!2,\!\dots\!,T\!-\!T_{g}^{D}\!\!+\!1$ (1f)
$u_{g,t}\in\{0,1\}\quad\forall g,t$ (1g)
$P_{h,t}^{shed}\geq 0\quad\forall h,t$ (1h)
In detail, during an operation horizon of $T$ periods, the decision variables are categorized into two groups. The first-stage decision variables are binary variables $u_{{g},{t}}$ determining the on/off states for the $g$-th unit in time period $t$. Since these first-stage decisions pertain to day-ahead scheduling, they are independent of realizations of uncertainties. Subsequently, when wait-and-see decisions are made, uncertainties in the power system are realized in terms of the output of wind farms on the supply side and the load on the demand side. The corresponding random variables are defined as $P_{{h},{t}}^{Wind}$ and $D_{{h},{t}}^{Load}$, respectively, for period $t$ for scenario $h$ in the scenario set $\Omega$ with probability $\pi\left(\xi_{h}\right)$. Due to uncertainties in both supply and demand in the second stage, load shedding $P_{{h},{t}}^{shed}$ is introduced to ensure the feasibility. Moreover, $C_{g}^{quad}$, $C_{g}^{prim}$ and $C_{g}^{cons}$ are cost coefficients. $C^{shed}$ is the penalty cost of load shedding. $P_{g}^{min}$ and $P_{g}^{max}$ donate the minimum and maximum power output. $P_{g}^{rd}$ and $P_{g}^{ru}$ are ramp-up and ramp-down rate limits. $T_{g}^{U}$ and $T_{g}^{D}$ denote the minimum up-time and down-time durations, respectively. The objective in SUC minimizes the total cost, consisting of constant production costs and the quadratic fuel costs of thermal units in (1a). At last, (1d) represents the power balance constraints for any given period in any scenario.
@@ II-A2 Solution Algorithm Based on Benders Decomposition
The BD method has been widely applied in large-scale power system optimization such as the SUC to reduce its computational burden[14, 23, 39]. It decouples complex, multi-constraint problems into small-scale subproblems that can be solved in parallel. In (1b), the first-stage variables are incorporated into the second-stage constraints, leading to a coupling between the two stages. To decouple the two stages, BD divides the SUC problem (1a)-(1h) into two sections, i.e., a mixed-integer master problem and independent scenario-based quadratic subproblems. The master problem captures the first-stage commitment decisions $u_{g,t}$. At iteration $k$, the master problem is written as,
$\min_{u_{g,t}^{k},\Upsilon^{k}}$ $\sum_{g=1}^{N}\sum_{t=1}^{T}C_{g}^{cons}u_{g,t}^{k}+\Upsilon^{k}$ (2a)
@@
$\hskip 56.9055pt\left.+C_{g}^{prim}P_{h,g,t}^{G,l}\bigg]+\sum_{t=1}^{T}C^{shed}P_{h,t}^{shed,l}\right\}$
@@
$\sum_{\varsigma=t}^{t+T_{g}^{U}-1}\!\!\!\!u_{g,\varsigma}\!\!\;\geq\;\!\!T_{g}^{U}\,\!\!\bigl(u_{g,t}\!-\!u_{g,t-1}\bigr)\!\!\!\quad\forall g,\;\!\!t\!=\!2,\!\dots\!,T\!-\!T_{g}^{U}\!\!+1\!\!\!$ (2d)
$\sum_{\varsigma=t}^{t+T_{g}^{D}-1}\!\!\!\!\!\bigl(1\!-\!u_{g,\varsigma}\bigr)\;\!\!\geq\;\!\!T_{g}^{D}\,\!\!\bigl(u_{g,t-1}\!-\!u_{g,t}\bigr)\!\!\!\!\quad\forall g,\;\!\!t\!=\!2,\!\dots\!,T\!-\!T_{g}^{D}\!\!+\!1$ (2e)
$u_{g,t}^{k}\in\{0,1\}\quad\forall g,t$ (2f)
where $\theta_{g,t}^{l}\left(\xi_{h}\right)$s are the optimal dual variables; $P_{h,g,t}^{G,l}$ and $P_{h,t}^{shed,l}$ are the optimal primal variables. Both are introduced by the subproblems. Note that, due to the introduction of a continuous Benders lower bound (LB), $\Upsilon$, the master problem is characterized as an MIP problem rather than a pure integer program.
The subproblem for scenario $h$ is given as (3a)-(3f) below. As each scenario is independent, the subproblem is decoupled and can be treated separately. The main objective is to strategize the optimal generation levels and corresponding load shedding for each scenario.
@@
In each subproblem, $u_{g,t}^{fixed}$ is a newly introduced auxiliary continuous variable and $u_{g,t}^{k}$ is the optimal decision variable of the master problem. As a result, (3) results in an upper bound (UB) of (1). Meanwhile, due to the convexity of (3), the dual variable $\theta_{g,t}^{l}\left(\xi_{h}\right)$s in (3e) defines the sensitivity of subproblem’s objective. After collecting all sensitivities, a new Benders cut (2c) can be added in the next iteration and the master problem (2) becomes a LB to (1). Moreover, as subproblem is a continuous quadratic programming, it can be efficiently solved on CPUs in parallel.
The BD method decouples the first-stage and second-stage operations, significantly reducing the problem size. It is noteworthy that all integer variables are now included in the master problem where QC holds the potential to further improve its computational efficiency.